OpenStudy (anonymous):

y^2-2y+1=9 give exact and approximate solutions to three decimal places

OpenStudy (anonymous):

\[y^2-2y+1-9=0\]\[y^2-2y-8=0\]\[(y+2)(y-4)=0\]y=-2 y=4